OpenStudy (anonymous):

please help me: Select the equation of a line that is parallel to the line on the graph and passes through the point (2, 3).

OpenStudy (anonymous):

Do you have a link or pic of it?

OpenStudy (anonymous):

i can try to get it

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

i can try to get it

OpenStudy (anonymous):

there it is

OpenStudy (anonymous):

ok so you want the equation right for the parallel line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = -1/2x + 8 :)

OpenStudy (anonymous):

y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = - 1 over 2 x + 4

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

then are you sure its point 2,3?